\begin{thebibliography}{99}

\bibitem{abram}
M. Abramowitz and I. A. Stegun (Eds), \textit{Handbook of Mathematical Functions with Formulas, Graphs, and Mathematical Tables}, National Bureau of Standards, Applied Mathematics Series \textbf{55}, 10th printing, Dover Publications, New York and Washington, 1972.

\bibitem{Allasia-Gior-Pecaric-MIA-02}
G. Allasia, C. Giordano, and J. Pe\v{c}ari\'c, \textit{Inequalities for the gamma function relating to asymptotic expasions}, Math. Inequal. Appl. \textbf{5} (2002), no.~3, 543\nobreakdash--555; available online at \url{https://doi.org/10.7153/mia-05-54}.

\bibitem{psi-alzer}
H. Alzer, \textit{On some inequalities for the gamma and psi functions}, Math. Comp. \textbf{66} (1997), no.~217, 373\nobreakdash--389; available online at \url{https://doi.org/10.1090/S0025-5718-97-00807-7}.

\bibitem{berg-pedersen-Alzer}
C. Berg and H. L. Pedersen, \textit{A completely monotonic function used in an inequality of Alzer}, Comput. Methods Funct. Theory \textbf{12} (2012), no.~1, 329\nobreakdash--341; available online at \url{https://doi.org/10.1007/BF03321830}.

\bibitem{Chen-Qi-Srivastava-09.tex}
C.-P. Chen, F. Qi, and H. M. Srivastava, \textit{Some properties of functions related to the gamma and psi functions}, Integral Transforms Spec. Funct. \textbf{21} (2010), no.~2, 153\nobreakdash--164; available online at \url{https://doi.org/10.1080/10652460903064216}.

\bibitem{Dubourdieu}
J. Dubourdieu, \textit{Sur un th\'{e}or\`eme de {M}. {S}. {B}ernstein relatif \`a la transformation de {L}aplace-{S}tieltjes}, Compositio Math. \textbf{7} (1939-40), 96\nobreakdash--111; available online at \url{http://www.numdam.org/item?id=CM_1940__7__96_0}. (French)

\bibitem{psi-proper-fraction-degree-two.tex}
B.-N. Guo and F. Qi, \textit{A completely monotonic function involving the tri-gamma function and with degree one}, Appl. Math. Comput. \textbf{218} (2012), no.~19, 9890\nobreakdash--9897; available online at \url{https://doi.org/10.1016/j.amc.2012.03.075}.

\bibitem{theta-new-proof.tex-BKMS}
B.-N. Guo and F. Qi, \textit{Two new proofs of the complete monotonicity of a function involving the psi function}, Bull. Korean Math. Soc. \textbf{47} (2010), no.~1, 103\nobreakdash--111; available online at \url{https://doi.org/10.4134/bkms.2010.47.1.103}.

\bibitem{Sharp-Ineq-Polygamma-Slovaca.tex}
B.-N. Guo, F. Qi, J.-L. Zhao, and Q.-M. Luo, \textit{Sharp inequalities for polygamma functions}, Math. Slovaca \textbf{65} (2015), no.~1, 103\nobreakdash--120; available online at \url{https://doi.org/10.1515/ms-2015-0010}.

\bibitem{Koumandos-jmaa-06}
S. Koumandos, \textit{Remarks on some completely monotonic functions}, J. Math. Anal. Appl. \textbf{324} (2006), no.~2, 1458\nobreakdash--1461; available online at \url{https://doi.org/10.1016/j.jmaa.2005.12.017}.

\bibitem{Koumandos-Pedersen-09-JMAA}
S. Koumandos and H. L. Pedersen, \textit{Completely monotonic functions of positive order and asymptotic expansions of the logarithm of Barnes double gamma function and Euler's gamma function}, J. Math. Anal. Appl. \textbf{355} (2009), no.~1, 33\nobreakdash--40; available online at \url{https://doi.org/10.1016/j.jmaa.2009.01.042}.

\bibitem{merkle-jmaa-96}
M. Merkle, \textit{Logarithmic convexity and inequalities for the gamma function}, J. Math. Anal. Appl. \textbf{203} (1996), no.~2, 369\nobreakdash--380; available online at \url{https://doi.org/10.1006/jmaa.1996.0385}.

\bibitem{mpf-1993}
D. S. Mitrinovi\'c, J. E. Pe\v{c}ari\'c, and A. M. Fink, \textit{Classical and New Inequalities in Analysis}, Kluwer Academic Publishers, Dordrecht-Boston-London, 1993; available online at \url{https://doi.org/10.1007/978-94-017-1043-5}.

\bibitem{Mortici-AA-10-134}
C. Mortici, \textit{Very accurate estimates of the polygamma functions}, Asympt. Anal. \textbf{68} (2010), no.~3, 125\nobreakdash--134; available online at \url{https://doi.org/10.3233/ASY-2010-0983}.

\bibitem{CAM-D-18-00067.tex}
F. Qi, \textit{A double inequality for the ratio of two non-zero neighbouring Bernoulli numbers}, J. Comput. Appl. Math. \textbf{351} (2019), 1\nobreakdash--5; available online at \url{https://doi.org/10.1016/j.cam.2018.10.049}.

\bibitem{AIMS-Math-2019595.tex}
F. Qi, \textit{Completely monotonic degree of a function involving trigamma and tetragamma functions}, AIMS Math. \textbf{5} (2020), no.~4, 3391\nobreakdash--3407; available online at \url{https://doi.org/10.3934/math.2020219}.

\bibitem{CM-degree-tri-p5.tex}
F. Qi, \textit{Completely monotonic degree of remainder of asymptotic expansion of trigamma function}, arXiv preprint (2020), available online at \url{https://arxiv.org/abs/2003.05300v1}.

\bibitem{Qi-Agar-Surv-JIA.tex}
F. Qi and R. P. Agarwal, \textit{On complete monotonicity for several classes of functions related to ratios of gamma functions}, J. Inequal. Appl. \textbf{2019}, Paper No.~36, 42~pages; available online at \url{https://doi.org/10.1186/s13660-019-1976-z}.

\bibitem{22ICFIDCAA-Filomat.tex}
F. Qi and W.-H. Li, \emph{Integral representations and properties of some functions involving the logarithmic function}, Filomat \textbf{30} (2016), no.~7, 1659\nobreakdash--1674; available online at \url{https://doi.org/10.2298/FIL1607659Q}.

\bibitem{CAM-D-18-02975.tex}
F. Qi and A.-Q. Liu, \textit{Completely monotonic degrees for a difference between the logarithmic and psi functions}, J. Comput. Appl. Math. \textbf{361} (2019), 366\nobreakdash--371; available online at \url{https://doi.org/10.1016/j.cam.2019.05.001}.

\bibitem{Mansour-Qi-CMD.tex-arXiv}
F. Qi and M. Mahmoud, \textit{Completely monotonic degrees of remainders of asymptotic expansions of the digamma function}, arXiv preprint (2019), available online at \url{https://arxiv.org/abs/1912.07989v2}.

\bibitem{Mansour-Qi-CMD.tex-HAL}
F. Qi and M. Mahmoud, \textit{Completely monotonic degrees of remainders of asymptotic expansions of the digamma function}, HAL preprint (2019), available online at \url{https://hal.archives-ouvertes.fr/hal-02415224v1}.

\bibitem{Schilling-Song-Vondracek-2nd}
R. L. Schilling, R. Song, and Z. Vondra\v{c}ek, \textit{Bernstein Functions}, 2nd ed., de Gruyter Studies in Mathematics \textbf{37}, Walter de Gruyter, Berlin, Germany, 2012; available online at \url{https://doi.org/10.1515/9783110269338}.

\bibitem{haerc1}
H. van Haeringen, \textit{Completely monotonic and related functions}, J. Math. Anal. Appl. \textbf{204} (1996), no.~2, 389\nobreakdash--408; available online at \url{https://doi.org/10.1006/jmaa.1996.0443}.

\bibitem{widder}
D. V. Widder, \textit{The Laplace Transform}, Princeton University Press, Princeton, 1946.

\bibitem{Xu=Cen-Qi-Conj-JIA2020}
A.-M. Xu and Z.-D. Cen, \textit{Qi's conjectures on completely monotonic degrees of remainders of asymptotic formulas of di- and tri-gamma functions}, J. Inequal. Appl. \textbf{2020}, Paper No.~83, 10~pages; available online at \url{https://doi.org/10.1186/s13660-020-02345-5}.

\bibitem{Zhao-Chu-JIA-392431}
T.-H. Zhao and Y.-M. Chu, \textit{A class of logarithmically completely monotonic functions associated with a gamma function}, J. Inequal. Appl. 2010, Art. ID~392431, 11~pages; available online at \url{https://doi.org/10.1155/2010/392431}.

\end{thebibliography}

\documentclass[reqno]{amsart}%
\usepackage{amssymb,setspace,mathtools}
\usepackage{ifpdf}
\ifpdf
 \usepackage[hyperindex]{hyperref}
\else
 \expandafter\ifx\csname dvipdfm\endcsname\relax
 \usepackage[hypertex,hyperindex]{hyperref}%
 \else
 \usepackage[dvipdfm,hyperindex]{hyperref}%
 \fi
\fi
\allowdisplaybreaks[4]
\numberwithin{equation}{section}
\theoremstyle{plain}
\newtheorem{thm}{Theorem}[section]
\theoremstyle{remark}
\newtheorem{rem}{Remark}[section]
\newcommand{\td}{\textup{d}}
\newcommand{\cmdeg}[1]{\sideset{}{_\textup{cm}^{#1}}\deg}

\begin{document}

\title[A double inequality for completely monotonic degree]
{A double inequality for completely monotonic degree of a remainder for an asymptotic expansion of the trigamma function}

\author[F. Qi]{Feng Qi}
\address{Institute of Mathematics, Henan Polytechnic University, Jiaozuo 454010, Henan, China\\
College of Mathematics and Physics, Inner Mongolia University for Nationalities, Tongliao 028043, Inner Mongolia, China\\
School of Mathematical Sciences, Tianjin Polytechnic University, Tianjin 300387, China}
\email{\href{mailto: F. Qi <qifeng618@gmail.com>}{qifeng618@gmail.com}, \href{mailto: F. Qi <qifeng618@hotmail.com>}{qifeng618@hotmail.com}, \href{mailto: F. Qi <qifeng618@qq.com>}{qifeng618@qq.com}}
\urladdr{\url{https://qifeng618.wordpress.com}}

\dedicatory{Dedicated to people facing and battling COVID-19}

\begin{abstract}
In the paper, the author presents a double inequality for completely monotonic degree of a remainder for an asymptotic expansion of the trigamma function. This result partially confirms one in a series of conjectures on completely monotonic degrees of remainders of asymptotic expansions for the logarithm of the gamma function and for polygamma functions.
\end{abstract}

\keywords{completely monotonic degree; remainder; asymptotic expansion; trigamma function; completely monotonic function; conjecture; gamma function; polygamma function}

\subjclass[2010]{Primary 33B15; Secondary 26A48, 26A51, 30E15, 44A10}

\thanks{This paper was typeset using\AmS-\LaTeX}

\maketitle
\tableofcontents

\section{Simple preliminaries}
In the literature~\cite[Section~6.4]{abram}, the functions
\begin{equation*}
\Gamma(w)=\int_0^{\infty}s^{w-1}e^{-s}\td s\quad\text{and}\quad
\psi(w)=[\ln\Gamma(w)]'=\frac{\Gamma'(w)}{\Gamma(w)}
\end{equation*}
for $\Re(w)>0$ are known as the classical Euler gamma function and the digamma function. Moreover, the functions $\psi'(w)$, $\psi''(w)$, $\psi^{(3)}(w)$, and $\psi^{(4)}(w)$ are known as the tri-, tetra-, penta-, and hexa-gamma functions respectively. As a whole, all the derivatives $\psi^{(k)}(w)$ for $k\ge0$ are known as the polygamma functions~\cite{berg-pedersen-Alzer, Zhao-Chu-JIA-392431}.
\par
Recall from the chapters~\cite[Chapter~XIII]{mpf-1993}, \cite[Chapter~1]{Schilling-Song-Vondracek-2nd}, and~\cite[Chapter~IV]{widder} that, if $\phi(t)$ defines on an interval $I$ and satisfies
\begin{equation}\label{cmf-dfn-ineq}
(-1)^{k}\phi^{(k)}(t)\geq0
\end{equation}
for all $t\in I$ and $k\ge0$, then it is called a completely monotonic function on $I$.
Theorem~12b on~\cite[p.~161]{widder} reads that $\phi(t)$ is completely monotonic on $(0,\infty)$ if and only if
\begin{equation}\label{Theorem12b-Laplace}
\phi(t)=\int_0^\infty e^{-ts}\td\sigma(s)
\end{equation}
converges for all $t\in(0,\infty)$, where $\sigma(s)$ is nondecreasing on $(0,\infty)$. The integral representation~\eqref{Theorem12b-Laplace} is equivalent to say that $\phi(t)$ is completely monotonic on $(0,\infty)$ if and only if it is a Laplace transform of $\sigma(s)$ on $(0,\infty)$.
A result in~\cite[p.~98]{Dubourdieu} and~\cite[p.~395]{haerc1}\label{CM-not=0} asserts that, unless $\phi(t)$ is a trivial completely monotonic function, that is, a nonnegative constant on $(0,\infty)$, those inequalities in~\eqref{cmf-dfn-ineq} are all strict on $(0,\infty)$. Why do we investigate completely monotonic functions? One can find historic answers in two monographs~\cite{Schilling-Song-Vondracek-2nd,widder} and closely related references therein.
\par
Let $\phi(t)$ be defined on $(0,\infty)$ and $\phi(\infty)=\lim_{x\to\infty}\phi(t)$. If $t^r[\phi(t)-\phi(\infty)]$ is completely monotonic on $(0,\infty)$, but $t^{r+\varepsilon}[\phi(t)-\phi(\infty)]$ is not for any number $\varepsilon>0$, then $r$ is called the completely monotonic degree of $\phi(t)$ with respect to $t\in(0,\infty)$; if $t^r[\phi(t)-\phi(\infty)]$ is completely monotonic on $(0,\infty)$ for all $r\in\mathbb{R}$, then the completely monotonic degree of $\phi(t)$ with respect to $t\in(0,\infty)$ is said to be $\infty$. The notation $\cmdeg{t}[\phi(t)]$ was designed in~\cite{psi-proper-fraction-degree-two.tex} to denote the completely monotonic degree $r$ of $\phi(t)$ with respect to $t\in(0,\infty)$. Why do we investigate completely monotonic degrees? One can find significant answers in the second paragraph of~\cite[Remark~1]{CAM-D-18-02975.tex} or in the papers~\cite{AIMS-Math-2019595.tex, Qi-Agar-Surv-JIA.tex, 22ICFIDCAA-Filomat.tex} and closely related references therein.

\section{Motivations}
It is well-known in~\cite[p.~257, 6.1.40]{abram} and~\cite[p.~260, 6.4.11]{abram} that
\begin{equation}\label{ln-gamma-symp-eq}
\ln\Gamma(w)\sim\biggl(w-\frac12\biggr)\ln w-w+\frac12\ln(2\pi)+\sum_{k=1}^\infty \frac{B_{2k}}{2k(2k-1)w^{2k-1}}
\end{equation}
and
\begin{equation}\label{asymptotic-polypsi}
\psi^{(n)}(w)\sim(-1)^{n-1}\biggl[\frac{(n-1)!}{w^n}+\frac{n!}{2w^{n+1}}+\sum_{k=1}^\infty B_{2k}\frac{(2k+n-1)!}{(2k)!w^{2k+n}}\biggr]
\end{equation}
as $w\to\infty$ in $\lvert \arg w\rvert<\pi$ for $n\ge0$, where an empty sum is understood to be $0$ and $B_{2k}$ for $k\ge1$ are known as the Bernoulli numbers which can be generated~\cite{CAM-D-18-00067.tex} by
\begin{equation*}
\frac{w}{e^w-1}=1-\frac{w}2+\sum_{k=1}^\infty B_{2k}\frac{w^{2k}}{(2k)!}, \quad| w| <2\pi.
\end{equation*}
Stimulated by the asymptotic expansions~\eqref{ln-gamma-symp-eq} and~\eqref{asymptotic-polypsi}, many mathematicians considered and investigated the remainders
\begin{equation*}
R_n(t)=(-1)^n\Biggl[\ln\Gamma(t)-\biggl(t-\frac12\biggr)\ln t+t-\frac12\ln(2\pi) -\sum_{k=1}^n\frac{B_{2k}}{2k(2k-1)}\frac1{t^{2k-1}}\Biggr]
\end{equation*}
and its derivatives, where $n\ge0$. For detailed information, please refer to the papers~\cite{Allasia-Gior-Pecaric-MIA-02, merkle-jmaa-96}, \cite[Theorem~8]{psi-alzer}, \cite[Section~1.4]{Sharp-Ineq-Polygamma-Slovaca.tex}, \cite[Theorem~2]{Koumandos-jmaa-06}, \cite[Theorem~2.1]{Koumandos-Pedersen-09-JMAA}, \cite[Theorem~3.1]{Mortici-AA-10-134}, and closely related references therein.
\par
In~\cite[Section~4]{CAM-D-18-02975.tex} and~\cite{Mansour-Qi-CMD.tex-arXiv, Mansour-Qi-CMD.tex-HAL}, the author posed, modified, and finalized the following conjectures:
\begin{enumerate}
\item
the completely monotonic degrees
\begin{align}
\cmdeg{t}\biggl[\ln t-\frac{1}{2t}-\psi(t)\biggr]&=1, \label{pis-ln-deg=1}\\
\cmdeg{t}\biggl[\frac{1}{t}+\frac{1}{2t^2}+\frac{1}{6t^3}-\psi'(t)\biggr]&=3, \label{pis-ln-deg=3}\\
\cmdeg{t}\biggl[\psi'(t)-\biggl(\frac{1}{t}+\frac{1}{2t^2}+\frac{1}{6t^3}-\frac{1}{30t^5}\biggr)\biggr]&=4 \label{pis-ln-deg=5}
\end{align}
with respect to $t\in(0,\infty)$ are valid;
\item
when $m=0$, the completely monotonic degrees of $R_n(t)$ with respect to $t\in(0,\infty)$ satisfy
\begin{equation}\label{R-01-0th-deg}
\cmdeg{t}[R_0(t)]=0,\quad \cmdeg{t}[R_1(t)]=1,
\end{equation}
and
\begin{equation}\label{R-n-0th-deg}
\cmdeg{t}[R_n(t)]=2(n-1), \quad n\ge2;
\end{equation}
\item
when $m=1$, the completely monotonic degrees of $-R_n'(t)$ with respect to $t\in(0,\infty)$ satisfy
\begin{equation}\label{R0p=1-R1p=2}
\cmdeg{t}[-R_0'(t)]=1,\quad \cmdeg{t}[-R_1'(t)]=2,
\end{equation}
and
\begin{equation}\label{Qi-Conj-Gam-Deg-der1(2n-1)}
\cmdeg{t}[-R_n'(t)]=2n-1, \quad n\ge2;
\end{equation}
\item
when $m\ge2$, the completely monotonic degrees of $(-1)^mR_n^{(m)}(t)$ with respect to $t\in(0,\infty)$ satisfy
\begin{equation}\label{R0mR1m-conj-Eq}
\cmdeg{t}\bigl[(-1)^mR_0^{(m)}(t)\bigr]=m, \quad \cmdeg{t}\bigl[(-1)^mR_1^{(m)}(t)\bigr]=m+1,
\end{equation}
and
\begin{equation}\label{R-n-m-th-deg}
\cmdeg{t}\bigl[(-1)^mR_n^{(m)}(t)\bigr]=m+2(n-1), \quad n\ge2.
\end{equation}
\end{enumerate}
\par
In~\cite[Theorem~1]{psi-alzer}, \cite[Theorem~1]{theta-new-proof.tex-BKMS}, \cite[Theorem~2.1]{Mansour-Qi-CMD.tex-arXiv}, \cite[Theorem~2.1]{Mansour-Qi-CMD.tex-HAL}, and~\cite[Theorem~3]{Xu=Cen-Qi-Conj-JIA2020}, the first conjecture in~\eqref{R0p=1-R1p=2}, which is equivalent to~\eqref{pis-ln-deg=1}, was unconsciously or consciously verified again and again.
\par
In~\cite[Theorem~2.1]{Koumandos-Pedersen-09-JMAA}, it was proved that
\begin{equation*}
\cmdeg{t}\bigl[R_n(t)\bigr]\ge n, \quad n\ge0.
\end{equation*}
This result is weaker than those conjectures in~\eqref{R-n-0th-deg}.
\par
In~\cite[Theorem~1]{Chen-Qi-Srivastava-09.tex}, \cite[Theorem~2]{CAM-D-18-02975.tex}, and~\cite[Theorem~4]{Xu=Cen-Qi-Conj-JIA2020}, the second conjecture in~\eqref{R0p=1-R1p=2} was proved once again.
\par
In~\cite[Theorem~2.1]{Mansour-Qi-CMD.tex-arXiv} and~\cite[Theorem~2.1]{Mansour-Qi-CMD.tex-HAL}, those five conjectures in~\eqref{R-01-0th-deg}, \eqref{R0p=1-R1p=2}, and~\eqref{Qi-Conj-Gam-Deg-der1(2n-1)} were confirmed.
\par
In~\cite[Theorems~1 and~2]{Xu=Cen-Qi-Conj-JIA2020}, it was acquired that
\begin{equation*}
\cmdeg{t}\bigl[(-1)^2R_0''(t)\bigr]=2 \quad\text{and}\quad
\cmdeg{t}\bigl[(-1)^2R_1''(t)\bigr]=3.
\end{equation*}
These two results confirm conjectures in~\eqref{R0mR1m-conj-Eq} just for the case $m=2$. The latter is equivalent to the conjecture~\eqref{pis-ln-deg=3}.
\par
The main aim of this paper is to partially confirm the conjecture~\eqref{pis-ln-deg=5}, a special case $m=n=2$ of the conjecture~\eqref{R-n-m-th-deg}, by the double inequality
\begin{equation}\label{Rn2m2deg=4}
4\le\cmdeg{t}\bigl[(-1)^2R_2''(t)\bigr]\le5.
\end{equation}

\section{A double inequality and its proof}
The conjecture~\eqref{pis-ln-deg=5}, or say, a special case $m=n=2$ of the conjecture in~\eqref{R-n-m-th-deg}, can be partially confirmed by the following theorem.

\begin{thm}\label{pis-ln-deg=5-thm}
The completely monotonic degree of the function
\begin{equation*}
Q(t)=\psi'(t)-\frac{1}{t}-\frac{1}{2t^2}-\frac{1}{6t^3}+\frac{1}{30t^5}
\end{equation*}
with respect to $t\in(0,\infty)$ is not less that $4$ and not greater than $5$. In other words, the double inequality~\eqref{Rn2m2deg=4} is valid.
\end{thm}

\begin{proof}
Making use of the integral representations
\begin{equation*}
\frac1{w^\tau}=\frac1{\Gamma(\tau)}\int_0^\infty s^{\tau-1}e^{-ws}\td s, \quad \Re(w), \Re(\tau)>0
\end{equation*}
and
\begin{equation*}
\psi^{(k)}(w)=(-1)^{k+1}\int_{0}^{\infty}\frac{s^{k}}{1-e^{-s}}e^{-ws}\td s, \quad \Re(w)>0, k\ge1
\end{equation*}
in~\cite[p.~260, 6.4.1]{abram} and~\cite[p.~255, 6.1.1]{abram} and directly computing, we obtain
\begin{align*}
Q(t)&=\int_{0}^{\infty}\biggl(\frac{s}{1-e^{-s}}-1-\frac{s}{2}-\frac{s^2}{12}+\frac{s^4}{6!}\biggr)e^{-ts}\td s\\
&\triangleq\int_{0}^{\infty}h(s)e^{-ts}\td s,\\
h'(s)&=\frac{s^3+\bigl(s^3-30s+90\bigr)e^{2s}-2s\bigl(s^2+60\bigr)e^s-30s-90}{180(e^s-1)^2},\\
h''(s)&=\frac{e^{3s}\bigl(s^2-10\bigr)+3\bigl(s^2+20s+30\bigr)e^s-3\bigl(s^2-20s+30\bigr)e^{2s}-s^2+10}{60(e^s-1)^3},\\
h^{(3)}(s)&=\frac{se^{4s}+(90-34s)e^{3s}-114se^{2s}-2(17s+45)e^s+s}{30(e^s-1)^4},\\
h^{(4)}(s)&=\frac{\begin{bmatrix}e^{5s}+5(6s-25)e^{4s}+10(33s-35)e^{3s}\\
+10(33s+35)e^{2s}+5(6s+25)e^s-1\end{bmatrix}}{30(e^s-1)^5}\\
&=\frac{1}{30(e^s-1)^5}\sum_{k=7}^{\infty}\frac{\begin{bmatrix}
5^k+(110k-350)3^k+5(3k-50)2^{2k-1}\\
+(165k+350)2^{k} +30k+125\end{bmatrix}}{k!}s^k\\
&=\frac{1}{60(e^s-1)^5}\biggl(\frac{5s^7}{7}+\frac{25s^8}{14}+\frac{193s^9}{84}+\frac{85s^{10}}{42} +\frac{5065s^{11}}{3696}+\dotsm\biggr)\\
&>0
\end{align*}
on $(0,\infty)$. Consecutively utilizing the L'H\^ospital rule gives
\begin{equation*}
\lim_{s\to0^+}h(s)=0,\quad\lim_{s\to0^+}h'(s)=0,\quad\lim_{s\to0^+}h''(s)=0,\quad\lim_{s\to0^+}h^{(3)}(s)=0.
\end{equation*}
Accordingly, consecutively integrating by parts results in
\begin{align*}
t^4Q(t)&=-t^3\int_{0}^{\infty}h(s)\frac{\td\bigl(e^{-ts}\bigr)}{\td s}\td s\\
&=-t^3\biggl[h(s)e^{-ts}\big|_{s=0^+}^{s=\infty}-\int_{0}^{\infty}h'(s)e^{-ts}\td s\biggr]\\
&=t^3\int_{0}^{\infty}h'(s)e^{-ts}\td s\\
&=-t^2\int_{0}^{\infty}h'(s)\frac{\td\bigl(e^{-ts}\bigr)}{\td s}\td s\\
&=-t^2\biggl[h'(s)e^{-ts}\big|_{s=0^+}^{s=\infty}-\int_{0}^{\infty}h''(s)e^{-ts}\td s\biggr]\\
&=t^2\int_{0}^{\infty}h''(s)e^{-ts}\td s\\
&=-t\int_{0}^{\infty}h''(s)\frac{\td\bigl(e^{-ts}\bigr)}{\td s}\td s\\
&=-t\biggl[h''(s)e^{-ts}\big|_{s=0^+}^{s=\infty}-\int_{0}^{\infty}h^{(3)}(s)e^{-ts}\td s\biggr]\\
&=t\int_{0}^{\infty}h^{(3)}(s)e^{-ts}\td s\\
&=-\int_{0}^{\infty}h^{(3)}(s)\frac{\td\bigl(e^{-ts}\bigr)}{\td s}\td s\\
&=-\biggl[h^{(3)}(s)e^{-ts}\big|_{s=0^+}^{s=\infty}-\int_{0}^{\infty}h^{(4)}(s)e^{-ts}\td s\biggr]\\
&=\int_{0}^{\infty}h^{(4)}(s)e^{-ts}\td s.
\end{align*}
Consequently, the function $t^4Q(t)$ is completely monotonic on $(0,\infty)$. This means that
\begin{equation}\label{Q(t)ge4}
\cmdeg{t}[Q(t)]\ge4.
\end{equation}
\par
For $\Re(w)>0$ and $k\ge1$, we have the recurrent relation
\begin{equation*}
\psi^{(k-1)}(w+1)=\psi^{(k-1)}(w)+(-1)^{k-1}\frac{(k-1)!}{w^k}.
\end{equation*}
See~\cite[p.~260, 6.4.6]{abram}. From this, it follows that
\begin{equation}
\begin{aligned}\label{t2zero-lim-polyg}
\lim_{t\to0^+}\bigl[t^{k}\psi^{(k-1)}(t)\bigr]
&=\lim_{t\to0^+}\biggl(t^{k}\biggl[\psi^{(k-1)}(t+1)-(-1)^{k-1}\frac{(k-1)!}{t^{k}}\biggr]\biggr)\\
&=(-1)^{k}(k-1)!
\end{aligned}
\end{equation}
for $k\ge1$.
If $t^{4+\varepsilon}Q(t)$ for $\varepsilon\in\mathbb{R}$ were completely monotonic on $(0,\infty)$, by virtue of the result mentioned on page~\pageref{CM-not=0} and proved in~\cite[p.~98]{Dubourdieu} and~\cite[p.~395]{haerc1}, we see that the first derivative of $t^{4+\varepsilon}Q(t)$ should be negative on $(0,\infty)$, that is,
\begin{align*}
\bigl[t^{4+\varepsilon}Q(t)\bigr]'&=(4+\varepsilon)t^{3+\varepsilon}Q(t)+t^{4+\varepsilon}Q'(t)\\
&=Q(t)t^{3+\varepsilon}\biggl[(4+\varepsilon)+t\frac{Q'(t)}{Q(t)}\biggr]
\end{align*}
should be negative on $(0,\infty)$. Therefore, we acquire
\begin{equation*}
\varepsilon<-4-t\frac{Q'(t)}{Q(t)}
=-4-t\frac{\psi''(t)-\frac{1}{6t^6}+\frac{1}{2t^4}+\frac{1}{t^3}+\frac{1}{t^2}} {\psi'(t)+\frac{1}{30t^5}-\frac{1}{6t^3}-\frac{1}{2t^2}-\frac{1}{t}}
\to1, \quad t\to0^+
\end{equation*}
on $(0,\infty)$, where we used the limit~\eqref{t2zero-lim-polyg} in the last step. Consequently, we arrive at
\begin{equation}\label{Q(t)<5}
\cmdeg{t}[Q(t)]\le5.
\end{equation}
\par
Combining~\eqref{Q(t)ge4} with~\eqref{Q(t)<5} leads to the inequality~\eqref{Rn2m2deg=4}. The proof of Theorem~\ref{pis-ln-deg=5-thm} is complete.
\end{proof}

\section{An open problem}

For a given point $(a,b)$ on the first quadrant $\{(x,y):x,y>0\}$, let $y=f(x)$ be a differentiable function on $[0,a]$ with $a>0$, $f(0)=0$, and $f(a)=b>0$. Denote $\tan\theta(x)=f'(x)$ and
\begin{equation*}
S_{a,b}(f)=\int_{0}^{a}\cos\theta(x)\td x=\int_{0}^{a}\cos\arctan f'(x)\td x=\int_{0}^{a}\frac{1}{\sqrt{1+[f'(x)]^2}\,}\td x.
\end{equation*}
What is the minimum
$$
\inf\bigl\{S_{a,b}(f):f(0)=0,f(a)=b,f(x)\in C^1([0,a])\bigl\}?
$$
What is the differentiable function $f_m(x)$ on $[0,a]$ such that $S_{a,b}(f_m)$ attains the above infimum?

\begin{rem}
This paper is a slightly revised version of the preprint~\cite{CM-degree-tri-p5.tex}.
\end{rem}

\end{document}